OpenStudy (anonymous):

How do I find E(p)??????? Everytime I try I end up with q in the equation... Shark Inc. has determined that demand for its newest netbook model is given by lnq−1lnp+0.004p=7, where q is the number of netbooks Shark can sell at a price of p dollars per unit. Shark has determined that this model is valid for prices p≥100.

OpenStudy (campbell_st):

well start by rewriting the equation \[\ln(q) = - \ln(p) - 0.004p + 7\] or \[\ln(q) = 7 - \ln(p) - 0.004p\] the opposite og a base e log is the exponential e so raise each term as a power of e the log/ exponential law is \[e^{\ln(a)} = a\] so then you have \[e^{\ln(q)} = e^7 - e^{0.004p} - e^{\ln(p)} \] which simplifies to \[q = e^7 - e^{0.004p} - p\]

OpenStudy (anonymous):

does it matter if u simplify before you derive it?

OpenStudy (campbell_st):

well q can't be simplified....

OpenStudy (anonymous):

no i know that, but I mean use the q= to derive rather then deriving from lnq=....

OpenStudy (campbell_st):

yes... use the equation of q to differentiate

OpenStudy (anonymous):

so dq/dp would be -0.004p e^(0.004p) -1 ?

OpenStudy (campbell_st):

that looks ok...

OpenStudy (campbell_st):

but you shouldn't have the 0.004p, it should be just 0.004

OpenStudy (anonymous):

oh right because the p would be a 1 right?

OpenStudy (anonymous):

so then E(p)= (dq/dp) (p/q) so i'm getting (-0.004(e^(0.004p))/(e^7- e^(0.004p)) but its saying i'm still wrong....

OpenStudy (campbell_st):

well aren't you missing a few things \[E(p) = \frac{(-0.004e^{0.004p} -1)\times p}{e^7 - e^{0.004p} - p}\]